# RiverScope: High-Resolution River Masking Dataset

### v1

@@ Abstract
Datasetshttps://github.com/cvl-umass/riverscope
Figure 1: RiverScope presents a global, high-resolution satellite image dataset focused on rivers using PlanetScope (PBC 2024) and co-registered with SWOT (Vinogradova et al. 2025), SWORD (Altenau et al. 2021), and Sentinel-2 (ESA 2022). To the left we show the distribution and splits of our expert-labeled dataset, covering various geographic and hydrological contexts.
@@ 1 Introduction
Global surface waters form the circulatory system of the climate, playing a vital role in transporting water and the materials it carries across the planet (Alsdorf, Rodr´ıguez, and Lettenmaier 2007; Cazenave et al. 2004). These dynamics are shaped not only by natural phenomena like floods and erosion but also by human activities like dam construction and irrigation (Yang et al. 2022; Feng et al. 2022). Understanding and mapping surface water dynamics is increasingly important for climate resilience and environmental sustainability due to their role in agriculture (Tian et al. 2015; Van der Molen, Breeuwsma, and Boers 1998), hydropower (Koralay, Kara, and Kezik 2018; Wasti et al. 2022), ecological services (Grizzetti et al. 2019; Zhao et al. 2003), urban planning (Ellis 2013; Bernhardt et al. 2008), and transportation (Opher and Friedler 2010).
@@
While the hydrology community has made progress in locating and mapping surface water (i.e., hydrography) (Isikdogan, Bovik, and Passalacqua 2019; Valman et al. 2024), characterizing water dynamics like extent and flow remains challenging. Satellite imagery is effective at capturing water extent, but properties such as flow are not directly observable (Feng et al. 2019; Gleason and Durand 2020). Stream gauges, which directly measure flow, are sparse and insufficient to provide a global view (Group et al. 2001; Gleason and Hamdan 2017).
@@
These limitations have driven the launch of the Surface Water and Ocean Topography (SWOT) mission (Biancamaria et al. 2016), which uses radar interferometry to measure surface water elevation with unprecedented precision (Vinogradova et al. 2025). However, river discharge (flow) is still inferred, a calculation that critically depends on accurate river width measurements (Bjerklie et al. 2018; Hagemann, Gleason, and Durand 2017). To support this, datasets like the SWOT River Database (SWORD) (Altenau et al. 2021) were developed, but they rely on width estimates from coarser Landsat imagery (Allen and Pavelsky 2018).
@@
Spatial resolution remains a key bottleneck, especially for small rivers and fine-scale changes (Filippucci et al. 2022; Guo-an et al. 2001). Sentinel-2 (ESA 2022) and Landsat (Earth Resources Observation and Science Center(2020) EROS) imagery (10-30m/pixel) often miss these features (Feng et al. 2022, 2019; Flores et al. 2024), limiting the accuracy of river models. This motivates the need for higher-resolution, expertly labeled datasets.
@@
We introduce RiverScope, a densely annotated, global-scale dataset of high-resolution (3 m/pixel) PlanetScope (PBC 2024) imagery of rivers and adjacent water bodies. The dataset covers 2,577 $km^{2}$, comprising 500x500-pixel images sampled across diverse geographic and hydrological contexts. Each image is co-located with Sentinel-2, SWOT, and SWORD data within a $\pm$12-hour window (see Figure 1). This unique alignment enables the comparison of sensor performance on hydrologically relevant tasks, quantifying trade-offs between accuracy and cost—a critical consideration for scaling monitoring systems at agencies like NASA or ESA. Surface water masks were created by hydrology and machine learning experts through 100 hours of manual annotation, ensuring high-quality ground truth data. We leverage this dataset to establish benchmarks using hydrologically relevant evaluation metrics.
@@
    Water segmentation. We evaluate diverse deep learning models on PlanetScope imagery, including ImageNet-pretrained models and remote sensing-specific methods. We show that lightweight input adapters using all the multispectral PlanetScope bands outperform RGB-only models. Our models outperform established baselines including NDWI (McFeeters 1996) and custom CNN architectures (Valman et al. 2024). PlanetScope’s fine spatial resolution enables detection of detailed river structures often missed by lower-resolution sensors like Sentinel (see Figure 2). Models trained on coarse imagery often struggle with precise boundary delineation, highlighting the potential of fusing complementary modalities in our dataset to improve performance of Sentinel estimates.
@@
    River width estimation. We introduce a benchmark for river width—a hydrologically meaningful metric poorly captured by pixel-level segmentation metrics like IoU. Our best model, trained on PlanetScope imagery, achieves a median width error of 7.2 meters, substantially outperforming the prior state-of-the-art of 30 meters (Valman et al. 2024). Landsat- and Sentinel-derived widths show median errors of 45.0 and 39.0 meters, respectively. Notably, we also provide the first evaluation of SWOT-derived widths, which exhibit a median error of 41.9 meters at “nodes” spaced every 200 meters along the river.
@@ 2 Related Work
Previous works have introduced large-scale datasets based on satellites with global coverage, such as Sentinel and Landsat (Astruc et al. 2025; Bastani et al. 2023; Claverie et al. 2018). For example, SatlasPretrain (Bastani et al. 2023) presents a diverse dataset of Sentinel and NAIP (USDA Farm Service Agency n.d.) imagery labeled for object detection, scene classification, and segmentation. EuroSAT (Helber et al. 2019) and BigEarthNet (Sumbul et al. 2019) provide Sentinel-based datasets for image classification, while NASA periodically releases segmentation labels through the Harmonized Landsat-Sentinel program (Jones 2019; Claverie et al. 2018), including water body mapping products (Daroya et al. 2025).
While Sentinel and Landsat offer broad spatial and temporal coverage, their resolutions (10 m for Sentinel, 30 m for Landsat) limit their utility for fine-grained hydrological analysis, such as tracking narrow river dynamics (Mullen et al. 2023; Tayer et al. 2023; Harlan et al. 2023; Kim et al. 2024). High-resolution datasets remain limited and are often focused on urban or industrial settings, relying on aerial imagery with sparse global coverage and infrequent revisit rates (Xia et al. 2018; Waqas Zamir et al. 2019; Christie et al. 2018; Xia et al. 2017; Cheng, Han, and Lu 2017).
The PlanetScope mission (PBC 2024) addresses these limitations by offering globally available, high-resolution (3 m per pixel), near-daily imagery. This has enabled recent work on surface water monitoring using PlanetScope data (Mullen et al. 2023; Tayer et al. 2023; Harlan et al. 2023; Kim et al. 2024; Valman et al. 2024; Flores et al. 2024). However, these efforts have largely focused on specific regions and do not publicly release annotated datasets.
To address this gap, we introduce RiverScope, a densely labeled, global-scale dataset of high-resolution PlanetScope imagery for water segmentation focused on rivers. While Valman et al. (2024) presents a similar large-scale water segmentation dataset, RiverScope extends this line of work by incorporating co-registered measurements from complementary sensors—including Sentinel (ESA 2022), Surface Water and Ocean Topography (SWOT) (Fu et al. 2024; Biancamaria et al. 2016), and SWOT River Database (SWORD) (Altenau et al. 2021)—and by openly releasing the annotated data.
@@
While early studies in satellite image modeling focused primarily on low-resolution data (Bastani et al. 2023; Jakubik et al. 2023; Manas et al. 2021; Daroya et al. 2025), a growing body of work now targets applications requiring high-resolution imagery (Wolters, Bastani, and Kembhavi 2023; Flores et al. 2024; Tayer et al. 2023). For instance, Flores et al. (2024) used PlanetScope imagery to segment headwater streams—small, often unnamed tributaries that require fine spatial details. Beyond stream segmentation (Harlan et al. 2023; Mullen et al. 2023; Tayer et al. 2023; Valman et al. 2024; Qayyum et al. 2020), PlanetScope has been applied to river discharge estimation (Wang and Vivoni 2022; Lamine et al. 2021; Feng et al. 2019), reservoir monitoring (Karaman 2022; Mansaray et al. 2021; Perin et al. 2021), and water quality assessment (Mansaray et al. 2021; Niroumand-Jadidi et al. 2020).
Building on prior work (Valman et al. 2024), we focus on water segmentation using PlanetScope, and further extend to estimating river widths and validating outputs of complementary sources such as Sentinel, SWOT, and SWORD. Despite increased interest in high-resolution applications, there has been limited exploration of pretraining strategies (e.g., ImageNet (Deng et al. 2009), CLIP (Radford et al. 2021), MoCov3 (Chen, Xie, and He 2021)) and segmentation architectures (Long, Shelhamer, and Darrell 2015; Chen et al. 2017; Ronneberger, Fischer, and Brox 2015; Ranftl, Bochkovskiy, and Koltun 2021) specifically tailored for this domain. Many prior approaches rely on randomly initialized convolutional models (Valman et al. 2024; Akiva et al. 2021) or simple thresholding methods (Feng et al. 2019; Cooley et al. 2017).
In this work, we systematically benchmark a range of segmentation architectures and pretraining methods for high-resolution river segmentation and width estimation. Across all models and training strategies, higher-resolution imagery delivers higher accuracy, highlighting the clear advantage of fine spatial detail for both river segmentation and width estimation.
@@ 3.1 Data Sources
Since Ka-band radar is less sensitive to clouds, SWOT provides large-scale, repeatable monitoring of rivers, lakes, and reservoirs wider than about 50 m worldwide, offering an alternative to traditional gauge-based systems. Previous studies (Yao et al. 2025; Pitcher et al. 2019) show SWOT can adequately resolve rivers, but mask quality is affected by factors such as vegetation, wetlands, and nearby urban areas. Our dataset provides the first systematic evaluation of river width estimates derived from SWOT on a global scale.
@@
is a global database of rivers designed to support SWOT river products. It integrates multiple data sources including river widths estimated from Landsat imagery (30 m/pixel resolution), river flow characteristics extracted from digital elevation maps (Yamazaki et al. 2019), and identified human-made obstructions along river paths. SWORD structures rivers using points (nodes) spaced approximately 200 m apart, connected by line segments (reaches) around 10 km long (see Figure 2). These nodes and reaches serve as reference points from which SWOT data are obtained. SWORD also provides river width estimates at the node level, which are part of our dataset.
@@
data from the European Space Agency (ESA) provides 10 m/pixel to 60 m/pixel multispectral imagery, depending on the spectral band (the red, blue, green, and near infrared (NIR) bands have 10 m/pixel resolution). We download tiles that correspond to the same location and approximate time as PlanetScope imagery, while selecting tiles with the least cloud cover. Although water segmentation labels are available from WorldCover (Van De Kerchove et al. 2021; Zanaga et al. 2022), these only provide data for 2020 and 2021 with a 10 m/pixel resolution. There are 13 spectral bands available for Sentinel imagery. We include all available spectral bands as part of our dataset.
@@
has a constellation of satellites collecting 3 m/pixel resolution image data with a revisit frequency of 1 day. Each image has 4 bands corresponding to the red, green, blue, and NIR bands. Images of size $\times500$ pixels were collected from 2023 to 2024 to be within the same time frame ($\pm$ 12 hours) as SWOT. River sites spanning a broad range of widths were selected using width estimates from other sensors (e.g., SWORD) and were limited to SWOT’s fast sampling orbit (i.e., the path where SWOT makes frequent measurements) to maximize the number of spatially intersecting samples. Figure 1 shows the distribution of the obtained 1,145 images from different geographic locations. Images with no cloud cover were selected to ensure optimal visibility of rivers. We manually label water pixels in all these images to cover an area of 2,577 $km^{2}$ (see § 3.2). We publicly release normalized versions (i.e., min-max normalized pixel values per image) of the multispectral data. Links to the original PlanetScope products with raw values are also available, and can be purchased directly from (PBC 2024). Models presented here are trained on the normalized data.
@@ 3.2 PlanetScope Data Labeling
Unlike Sentinel images, PlanetScope images provide higher temporal and spatial resolution, enabling more detailed fine-grained image analyses. Consequently, these higher-resolution images were selected for manual annotation. Each PlanetScope image was manually annotated by one of 15 hydrology and river experts using a multi-scale pixel-wise annotator (Tangseng, Wu, and Yamaguchi 2017). Annotators were provided with the RGB image of the river to label, corresponding SWORD reaches and nodes overlaid on the images, a zoomed-out contextual view surrounding the target area, and a Google Maps link of the location. These supplementary information were included to enhance context and accuracy in labeling. Annotators also had interactive abilities such as zooming in and out of the target area, and the option of using a variety of labeling tools. Labels can be provided using a paintbrush, a polygon selector, and a superpixel selection tool that clusters similar pixels together. These interactive features streamlined the annotation process, ensuring comprehensive and efficient label coverage. A detailed description and visualization of the tool is provided in § A.2 (Appendix).
@@ 3.4 Tasks
For each node defined in SWORD, the ground truth river width is derived from PlanetScope water segmentation labels (see Figure 2). Given a satellite image, the corresponding river reach that spatially intersects the image is retrieved from SWORD. For each node along the river reach, the local slope is estimated, and a perpendicular (orthogonal) line is computed. River widths are then calculated by counting the number of water pixels along this orthogonal line and multiplying by the image resolution (3 m/pixel for PlanetScope) (Yang et al. 2019). These measurements serve as the ground truth river widths for SWORD nodes and provide the reference for evaluating other methods.
@@ 4.1 Training Details
We evaluate 27 models on RiverScope tasks based on varying segmentation models, backbones, and pretraining methods (Iakubovskii 2019) to see the effect of training with our dataset across different settings. We experiment on four semantic segmentation models covering different ways of handling features and multi-scale information: FPN (Long, Shelhamer, and Darrell 2015), DeepLabv3 (Chen et al. 2017), UNet (Ronneberger, Fischer, and Brox 2015), DPT (Ranftl, Bochkovskiy, and Koltun 2021). For each segmentation model, we then experiment with different backbones and available pretrained weights which are the basis of extracting features for each segmentation model. FPN, DeepLabv3, and UNet are applied to CNN and Swin-based backbones, while DPT is applied to ViT-based backbones. ResNet50 (RN50) (He et al. 2016), MobileNetv2 (MV2) (Sandler et al. 2018), Swin-T (Liu et al. 2021), Swin-B (Liu et al. 2021), ViT-B/16 (Dosovitskiy et al. 2021), and ViT-L/16 (Dosovitskiy et al. 2021) backbones are used. Pretraining methods also vary from supervised methods using SatlasNet (Bastani et al. 2023) and ImagetNet1k (Deng et al. 2009), and self-supervised methods using SeCo (Manas et al. 2021), MoCov3 (Chen, Xie, and He 2021), CLIP (Radford et al. 2021), Prithvi (Jakubik et al. 2023), and DINO (Caron et al. 2021). Among these, SeCo, SatlasNet, and Prithvi use satellite images for pretraining. Each of these configurations serves as the starting point before further fine-tuning with RiverScope.
All models are trained on the PlanetScope labeled images from RiverScope. Binary cross entropy loss (Eq. 1) is used to train all models. A predicted segmentation mask $\hat{\mathbf{Y}}_{i}\in\mathbb{R}^{W\times H}$ with pixels $\hat{y}_{i,(j,k)}$ is compared with the ground truth mask $\mathbf{Y}_{i}\in\mathbb{R}^{W\times H}$ using Eq. 1. The Adam optimizer (Kingma 2014) is used for training with the learning rate chosen from $10^{-1}$ to $10^{-6}$ based on the validation set performance.

### v2

@@ Abstract
Datasethttps://github.com/cvl-umass/riverscope
Figure 1: RiverScope presents a global, high-resolution satellite image dataset focused on rivers using PlanetScope (planetlabs) and co-registered with SWOT (vinogradova2025new), SWORD (altenau2021surface), and Sentinel-2 (esa2022sentinel). To the left we show the distribution and splits of our expert-labeled dataset, covering various geographic and hydrological contexts.
@@ 1 Introduction
Global surface waters form the circulatory system of the climate, playing a vital role in transporting water and the materials it carries across the planet (alsdorf2007measuring). These dynamics are shaped not only by natural phenomena like floods and erosion but also by human activities like dam construction and irrigation (yang2022impact). Understanding and mapping surface water dynamics is increasingly important for climate resilience and environmental sustainability due to their role in agriculture (tian2015modeling), hydropower (wasti2022climate), ecological services (zhao2003ecosystem), urban planning (ellis2013sustainable), and transportation (opher2010factors).
@@
While the hydrology community has made progress in locating and mapping surface water (i.e., hydrography) (valman2024ai), characterizing water dynamics like extent and flow remains challenging. Satellite imagery is effective at capturing water extent, but properties such as flow are not directly observable (feng2019comparing). Stream gauges, which directly measure flow, are sparse and insufficient to provide a global view (gleason2017crossing).
@@
These limitations have driven the launch of the Surface Water and Ocean Topography (SWOT) mission (biancamaria2016remote), which uses radar interferometry to measure surface water elevation with unprecedented precision (vinogradova2025new). However, river discharge (flow) is still inferred, a calculation that critically depends on accurate river width measurements (bjerklie2018satellite). To support this, datasets like the SWOT River Database (SWORD) (altenau2021surface) were developed, but they rely on width estimates from coarser Landsat imagery (allen2018global).
@@
Spatial resolution remains a key bottleneck, especially for small rivers and fine-scale changes (filippucci2022sentinel). Sentinel-2 (esa2022sentinel) and Landsat (eros2020_landsat_l2_c2) imagery (10-30m/pixel) often miss these features (flores2024mapping), limiting the accuracy of river models. This motivates the need for higher-resolution, expertly labeled datasets.
@@
We introduce RiverScope, a densely annotated, global-scale dataset of high-resolution (3 m/pixel) PlanetScope (planetlabs) imagery of rivers and adjacent water bodies. The dataset covers 2,577 $km^{2}$, comprising 500x500-pixel images sampled across diverse geographic and hydrological contexts. Each image is co-located with Sentinel-2, SWOT, and SWORD data within a $\pm$12-hour window (see Figure 1). This unique alignment enables the comparison of sensor performance on hydrologically relevant tasks, quantifying trade-offs between accuracy and cost—a critical consideration for scaling monitoring systems at agencies like NASA or ESA. Surface water masks were created by hydrology and machine learning experts through 100 hours of manual annotation, ensuring high-quality ground truth data. We leverage this dataset to establish benchmarks using hydrologically relevant evaluation metrics.
@@
    Water segmentation. We evaluate diverse deep learning models on PlanetScope imagery, including ImageNet-pretrained models and remote sensing-specific methods. We show that lightweight input adapters using all the multispectral PlanetScope bands outperform RGB-only models. Our models outperform established baselines including NDWI (mcfeeters1996use) and custom CNN architectures (valman2024ai). PlanetScope’s fine spatial resolution enables detection of detailed river structures often missed by lower-resolution sensors like Sentinel (see Figure 2). Models trained on coarse imagery often struggle with precise boundary delineation, highlighting the potential of fusing complementary modalities in our dataset to improve performance of Sentinel estimates.
@@
    River width estimation. We introduce a benchmark for river width—a hydrologically meaningful metric poorly captured by pixel-level segmentation metrics like IoU. Our best model, trained on PlanetScope imagery, achieves a median width error of 7.2 meters, substantially outperforming the prior state-of-the-art of 30 meters (valman2024ai). Landsat- and Sentinel-derived widths show median errors of 45.0 and 39.0 meters, respectively. Notably, we also provide the first evaluation of SWOT-derived widths, which exhibit a median error of 41.9 meters at “nodes” spaced every 200 meters along the river.
@@ 2 Related Work
Previous works have introduced large-scale datasets based on satellites with global coverage, such as Sentinel and Landsat (bastani2023satlaspretrain; claverie2018harmonized). For example, SatlasPretrain (bastani2023satlaspretrain) presents a diverse dataset of Sentinel and NAIP (usda_naip_imagery) imagery labeled for object detection, scene classification, and segmentation. EuroSAT (helber2019eurosat) and BigEarthNet (sumbul2019bigearthnet) provide Sentinel-based datasets for image classification, while NASA periodically releases segmentation labels through the Harmonized Landsat-Sentinel program (jones2019improved), including water body mapping products (daroya2025improving).
While Sentinel and Landsat offer broad spatial and temporal coverage, their resolutions (10 m for Sentinel, 30 m for Landsat) limit their utility for fine-grained hydrological analysis, such as tracking narrow river dynamics. High-resolution datasets remain limited and are often focused on urban or industrial settings, relying on aerial imagery with sparse global coverage and infrequent revisit rates (christie2018functional_fmow).
The PlanetScope mission (planetlabs) addresses these limitations by offering globally available, high-resolution (3 m per pixel), near-daily imagery. This has enabled recent work on surface water monitoring using PlanetScope data (valman2024ai; flores2024mapping). However, these efforts have largely focused on specific regions and do not publicly release annotated datasets.
To address this gap, we introduce RiverScope, a densely labeled, global-scale dataset of high-resolution PlanetScope imagery for water segmentation focused on rivers. While valman2024ai presents a similar large-scale water segmentation dataset, RiverScope extends this line of work by incorporating co-registered measurements from complementary sensors—including Sentinel (esa2022sentinel), Surface Water and Ocean Topography (SWOT) (biancamaria2016remote), and SWOT River Database (SWORD) (altenau2021surface)—and by openly releasing the annotated data.
@@
While early studies in satellite image modeling focused primarily on low-resolution data (bastani2023satlaspretrain; manas2021seasonal_seco; daroya2025improving), a growing body of work now targets applications requiring high-resolution imagery (flores2024mapping). For instance, flores2024mapping used PlanetScope imagery to segment headwater streams—small, often unnamed tributaries that require fine spatial details. Beyond stream segmentation (valman2024ai), PlanetScope has been applied to river discharge estimation, reservoir monitoring, and water quality assessment (wang2022mapping; flores2024mapping).
Building on prior work (valman2024ai), we focus on water segmentation using PlanetScope, and further extend to estimating river widths and validating outputs of complementary sources such as Sentinel, SWOT, and SWORD. Despite increased interest in high-resolution applications, there has been limited exploration of pretraining strategies (e.g., ImageNet (deng2009imagenet), CLIP (radford2021learning_clip), MoCov3 (chen2021empirical_mocov3)) and segmentation architectures (ronneberger2015u; ranftl2021vision) specifically tailored for this domain. Many prior approaches rely on randomly initialized convolutional models (valman2024ai) or simple thresholding methods (mcfeeters1996use).
In this work, we benchmark a range of segmentation architectures and pretraining methods for high-resolution river segmentation and width estimation. Across all models and training strategies, higher-resolution imagery delivers higher accuracy, highlighting the clear advantage of fine spatial detail for both river segmentation and width estimation.
@@ 3.1 Data Sources
Since Ka-band radar is less sensitive to clouds, SWOT provides large-scale, repeatable monitoring of rivers, lakes, and reservoirs wider than about 50 m worldwide, offering an alternative to traditional gauge-based systems. Previous studies (yao2025swot) show SWOT can adequately resolve rivers, but mask quality is affected by factors such as vegetation, wetlands, and nearby urban areas. Our dataset provides the first systematic evaluation of river width estimates derived from SWOT on a global scale.
@@
is a global database of rivers designed to support SWOT river products. It integrates multiple data sources including river widths estimated from Landsat imagery (30 m/pixel resolution), river flow characteristics extracted from digital elevation maps (yamazaki2019merit), and identified human-made obstructions along river paths. SWORD structures rivers using points (nodes) spaced approximately 200 m apart, connected by line segments (reaches) around 10 km long (see Figure 2). These nodes and reaches serve as reference points from which SWOT data are obtained. SWORD also provides river width estimates at the node level, which are part of our dataset.
@@
data from the European Space Agency (ESA) provides 10 m/pixel to 60 m/pixel multispectral imagery, depending on the spectral band (the red, blue, green, and near infrared (NIR) bands have 10 m/pixel resolution). We download tiles that correspond to the same location and approximate time as PlanetScope imagery, while selecting tiles with the least cloud cover. Although water segmentation labels are available from WorldCover (van2021esa_worldcover), these only provide data for 2020 and 2021 with a 10 m/pixel resolution. There are 13 spectral bands available for Sentinel imagery. We include all available spectral bands as part of our dataset.
@@
has a constellation of satellites collecting 3 m/pixel resolution image data with a revisit frequency of 1 day. Each image has 4 bands corresponding to the red, green, blue, and NIR bands. Images of size $\times500$ pixels were collected from 2023 to 2024 to be within the same time frame ($\pm$ 12 hours) as SWOT. River sites spanning a broad range of widths were selected using width estimates from other sensors (e.g., SWORD) and were limited to SWOT’s fast sampling orbit (i.e., the path where SWOT makes frequent measurements) to maximize the number of spatially intersecting samples. Figure 1 shows the distribution of the obtained 1,145 images from different geographic locations. Images with no cloud cover were selected to ensure optimal visibility of rivers. We manually label water pixels in all these images to cover an area of 2,577 $km^{2}$ (see § 3.2). We publicly release normalized versions (i.e., min-max normalized pixel values per image) of the multispectral data. Links to the original PlanetScope products with raw values are also available, and can be purchased directly from (planetlabs). Models presented here are trained on the normalized data.
@@ 3.2 PlanetScope Data Labeling
Unlike Sentinel images, PlanetScope images provide higher temporal and spatial resolution, enabling more detailed fine-grained image analyses. Consequently, these higher-resolution images were selected for manual annotation. Each PlanetScope image was manually annotated by one of 15 hydrology and river experts using a multi-scale pixel-wise annotator (tangseng2017looking). Annotators were provided with the RGB image of the river to label, corresponding SWORD reaches and nodes overlaid on the images, a zoomed-out contextual view surrounding the target area, and a Google Maps link of the location. These supplementary information were included to enhance context and accuracy in labeling. Annotators also had interactive abilities such as zooming in and out of the target area, and the option of using a variety of labeling tools. Labels can be provided using a paintbrush, a polygon selector, and a superpixel selection tool that clusters similar pixels together. These interactive features streamlined the annotation process, ensuring comprehensive and efficient label coverage. A detailed description and visualization of the tool is provided in § A.3 (Appendix).
@@ 3.4 Tasks
For each node defined in SWORD, the ground truth river width is derived from PlanetScope water segmentation labels (see Figure 2). Given a satellite image, the corresponding river reach that spatially intersects the image is retrieved from SWORD. For each node along the river reach, the local slope is estimated, and a perpendicular (orthogonal) line is computed. River widths are then calculated by counting the number of water pixels along this orthogonal line and multiplying by the image resolution (3 m/pixel for PlanetScope) (yang2019rivwidthcloud). These measurements serve as the ground truth river widths for SWORD nodes and provide the reference for evaluating other methods.
@@ 4.1 Training Details
We evaluate 27 models on RiverScope tasks based on varying segmentation models, backbones, and pretraining methods (Iakubovskii:2019) to see the effect of training with our dataset across different settings. We experiment on four semantic segmentation models covering different ways of handling features and multi-scale information: FPN (long2015fully), DeepLabv3 (chen2017deeplab), UNet (ronneberger2015u), DPT (ranftl2021vision). For each segmentation model, we then experiment with different backbones and available pretrained weights which are the basis of extracting features for each segmentation model. FPN, DeepLabv3, and UNet are applied to CNN and Swin-based backbones, while DPT is applied to ViT-based backbones. ResNet50 (RN50) (heDeepResidualLearning2016), MobileNetv2 (MV2) (sandler2018mobilenetv2), Swin-T (liu2021swin), Swin-B (liu2021swin), ViT-B/16 (dosovitskiy2020vit), and ViT-L/16 (dosovitskiy2020vit) backbones are used. Pretraining methods also vary from supervised methods using SatlasNet (bastani2023satlaspretrain) and ImagetNet1k (deng2009imagenet), and self-supervised methods using SeCo (manas2021seasonal_seco), MoCov3 (chen2021empirical_mocov3), CLIP (radford2021learning_clip), Prithvi (jakubik2310foundation_prithvi), and DINO (caron2021emerging). Among these, SeCo, SatlasNet, and Prithvi use satellite images for pretraining. Each of these configurations serves as the starting point before further fine-tuning with RiverScope.
All models are trained on the PlanetScope labeled images from RiverScope. Binary cross entropy loss (Eq. 1) is used to train all models. A predicted segmentation mask $\hat{\mathbf{Y}}_{i}\in\mathbb{R}^{W\times H}$ with pixels $\hat{y}_{i,(j,k)}$ is compared with the ground truth mask $\mathbf{Y}_{i}\in\mathbb{R}^{W\times H}$ using Eq. 1. The Adam optimizer is used for training with the learning rate chosen from $10^{-1}$ to $10^{-6}$ based on the validation set performance.